OpenStudy (anonymous):

verify the identity. justify each step. tanx+cotx=1/sinxcosx help im so confused on this problem i have no idea

OpenStudy (luigi0210):

rewrite the tanx and cotx

OpenStudy (anonymous):

Ok so to do this one you have to break it up into its components and factor them out

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

tanx= sinx/cosx cotx+= cosx/sinx

OpenStudy (anonymous):

ok i have that but where do i go from there?

OpenStudy (anonymous):

so both sides have a sinx and a cosx so u can multiply both sidesby it

OpenStudy (luigi0210):

\[tanx+cotx=\frac{ 1 }{ sinxcosx }\] \[\frac{ sinx }{ cosx }+\frac{ cosx }{ sinx }\] \[\frac{ sinx }{ cosx }*\frac{ sinx }{ sinx }+\frac{ cosx }{ sinx }*\frac{ cosx }{ cosx }\] \[\frac{ \sin^2x }{ sinxcosx }+\frac{ \cos^2x }{ sinxcosx }\] \[\frac{ \sin^2x+\cos^2x }{ sinxcosx }\] \[\frac{ 1 }{ sinxcosx }=\frac{ 1 }{ sinxcosx }\]

OpenStudy (luigi0210):

1. convert tanx and cotx into sin's and cos's 2. Get common denominators:sinxcosx 3. Add the terms up 4. Use trig identities and change sin^2x+cos^2x into 1

OpenStudy (anonymous):

thank you very much :)

OpenStudy (luigi0210):

Ha, you're welcome :)